OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} x \sin \frac{1}{x}\] I need to show all my work, but I don't know how to show it. I know sin(1/x) is always between 0 and 1, but sin(1/x) is getting close to 0 as x approaches infinity, and so I'm getting infinity * 0 which doesn't really make any sense.

OpenStudy (anonymous):

Introduce new variable\[u = \frac{1}{x}\]

OpenStudy (anonymous):

Right, thank you

OpenStudy (anonymous):

then u have\[\lim_{u\rightarrow 0}\frac{\sin u}{u}\]

OpenStudy (anonymous):

which is as well known = 1